# Lifelong and Universal Machine Learning Potentials for Chemical Reaction Network Explorations

### v1

@@ Abstract
Recent developments in computational chemistry facilitate the automated quantum chemical exploration of chemical reaction networks for the in-silico prediction of synthesis pathways, yield, and selectivity. However, the underlying quantum chemical energy calculations require vast computational resources, limiting these explorations severely in practice. Machine learning potentials (MLPs) offer a solution to increase computational efficiency, while retaining the accuracy of reliable first-principles data used for their training. Unfortunately, MLPs will be limited in their generalization ability within chemical (reaction) space, if the underlying training data is not representative for a given application. Within the framework of automated reaction network exploration, where new reactants or reagents composed of any elements from the periodic table can be introduced, this lack of generalizability will be the rule rather than the exception. Here, we therefore study the benefits and drawbacks of two MLP concepts in this context. Whereas universal MLPs are designed to cover most of the relevant chemical space in their training, lifelong MLPs push their adaptability by efficient continual learning of additional data. While the accuracy of the universal MLPs turns out to be not yet sufficient for reaction search trials without any fine-tuning, lifelong MLPs can reach chemical accuracy. We propose an improved learning algorithm for lifelong adaptive data selection yielding efficient integration of new data while previous expertise is preserved.
Lifelong Machine Learning Potentials, Universal Machine Learning Potentials, Chemical Reaction Networks, Continual Resilient (CoRe) Optimizer, Lifelong Adaptive Data Selection
@@ 1 Introduction
Accordingly, fast quantum chemical approaches are needed. However, they are plagued by drastic approximations that compromise not only energies but also molecular structures [32, 33]. As a consequence, a CRN constructed with an approximate quantum chemical approach (such as tight-binding density functional or semi-empirical theories) does not necessarily represent the network that would be obtained with a more reliable approach (such as density functional theory (DFT)). Hence, node and connection fidelity can be compromised. Local refinement by applying more accurate methods for the a-posteriori reoptimization of stable intermediates and transition states could improve molecular structures and, most importantly, their relative energies [34, 35, 36]. However, this requires a CRN to be qualitatively correct. Hence, the stationary points optimized with a fast approach must be complete as must be their connections by elementary reaction steps. It is therefore key to devise first-principles-based methods that are (I) fast to allow for efficient reactive event screening and that have (II) high structural fidelity so that the topology and structure of a CRN is not compromised. Hence, machine learning potentials (MLPs) hold the greatest promise to yield both, efficiency and structure fidelity [37, 38, 39, 40, 41, 42, 43].
@@
However, the requirements for MLPs in CRN exploration are challenging. (1) They must be quickly initializable and (2) they must not come with an unbearable overhead for their training. It is precisely the the goal of universal MLPs (uMLPs) to deliver a parametrization that is as general as possible to overcome the initialization issue and to avoid training at runtime [44, 45, 46]. Unfortunately, an approach without further fine-tuning on some given chemical system can suffer from insufficient accuracy. In order to tackle this challenge, we therefore proposed the concept of lifelong machine learning potentials (lMLP) to alleviate the overhead of training [47]. In contrast to conventionally trained MLPs, lMLPs can continuously adapt to additional data without forgetting previous knowledge, while they do not require training again on all previous data. Hence, they can offer high accuracy for some parts of chemical space studied, but they may still require additional quantum chemical calculations during their application.
@@
In this work, we assess the reliability, accuracy, and efficiency of both, uMLPs and lMLPs, for CRN explorations. Section 2 summarizes the concepts of lMLPs and uMLPs and introduces an improved algorithm for lifelong adaptive data selection. The computational details are compiled in Section 3. Section 4 starts with a presentation of the $\text{HCN}+\text{H}_{2}\text{O}$ CRN studied in this work. The section continues with a performance assessment of uMLPs compared to DFT reference data. Afterwards, continual learning of an lMLP on the CRN data is analyzed and compared to conventional iterative learning as well as to uMLP and DFT results.
@@ 2.1 Lifelong Machine Learning Potentials
Figure 1: (a) In iterative learning, a new machine learning model is trained from scratch on all data to integrate additional training data. (a) In lifelong learning, the training of a model is continued employing added data and only a subset of previous data to prevent forgetting. In this way, previously acquired knowledge of the model is exploited. We note that lifelong learning is here only based on rehearsal of data, while it can also exploit model parameter regularization and model architecture.
@@
With the help of continual learning, a closed and efficient loop in the evolution and application of an lMLP can be obtained (Figure 2). The process starts with some initial reference data. These data can be represented through universal descriptors such as element-embracing atom-centered symmetry functions (eeACSFs) [47], so that they can be learned by an lMLP utilizing, for example, the high-dimensional neural network potential (HDNNP) method [56, 57, 58] (see Supporting Information Section S1.1 of this work for an overview). We note that the lMLP concept can also be employed with other methods, including approaches which combine descriptor and potential into a single learnable representation. In any case, the representation must be able to handle any chemical structure to yield a generally applicable lMLP. For example, many MLP descriptors cannot efficiently represent a system with many different chemical elements because their descriptor vector size increases with an increasing number of different elements. As a consequence, the application of conventional atom-centered symmetry functions [59] is often restricted to systems with at most four different elements because of computational costs. By combination of (molecular) structure information with element information of the periodic table, eeACSFs can overcome this limitation [60, 47] (see Supporting Information Section S1.2 of this work for an overview). Moreover, in Supporting Information Section S1.2 of this work we propose alternative functions based on the bump function to represent the radial and angular structure within eeACSFs and we introduce the cube root-scaled-shifted (crss) scaling function for eeACSFs.
@@
To be able to deal with molecules with different total charges or spin multiplicities, individual lMLPs for each state can be trained. Alternatively, the second-generation HDNNP base model can be replaced by other models that can handle multiple states, such as fourth-generation HDNNPs [61]. Moreover, continual learning can also be applied to train atomic-property machine learning models (e.g., for atomic charges and spins) to calculate these efficiently, e.g., during molecular dynamics simulations.
@@
The trained model can subsequently be applied, whereby uncertainty quantification can be exploited to probe a pre-defined accuracy on the fly and to identify the need for new data points. In fact, uncertainty quantification is necessary to enable application of an lMLP at every training stage because it provides the confidence interval for analysis of the results. Uncertainties can be obtained, for example, by an ensemble or committee model [62, 63, 64, 65]. This can provide a quantitative estimate of small errors, while large errors can only be flagged. However, an indication of large errors is sufficient because only the need for additional data must be revealed. Those are then produced and fed into the continual training process until all errors are found to be sufficiently small.
@@ 2.2 Lifelong Adaptive Data Selection
The main ideas behind lADS are that, on the one hand, training data will be redundant if they are seldom trained in continual learning but still well represented. On the other hand, it is likely that data will be inconsistent with the majority of data if they are very often trained but still poorly represented. Consequently, lADS requires only a fraction of all training structures in each training step, applying a biased random selection. The probability for training of a given structure thereby needs to be adjusted according to how well the structure is represented by the model and how stable the quality of this representation is. In this way, both aforementioned ideas can be exploited. Moreover, learning can also focus on the insufficiently represented training data.
@@
with the loss functions $\mathcal{L}$ of the energy $E^{r}$ and atomic force components $F_{\alpha,n}^{r}$ of the $N_{\mathrm{atom}}^{r}$ atoms $n$ of each structure $r$,
@@
The main difference of the lADS algorithm presented in this work compared to its original implementation in Reference [47] is (I) the separate consideration of energy and force loss contributions. (II) the advanced algorithm adjusts the number of fitted structures per training step. (III) it employs an improved scheme to calculate the training probabilities and (IV) to update the selection determining properties. (V) it includes a maximum number of structures that can be classified as redundant per step and (VI) adds a scheme for fast integration of additional training data. In addition, (VII) it enables removing loss gradient contributions of inconsistent training data in the history of the CoRe optimizer.
@@
$L_{\mathrm{old}}^{\mathrm{max},E/F}\leftarrow\mathrm{max}\left(\mathbf{L}_{\mathrm{old}}^{\backslash\mathrm{NaN},S_{\mathrm{hist}}^{r,E}>0,E/F}\right)$
@@
$\absentfit\mathbf{P}_{\mathrm{good}}\leftarrow\mathbf{S}_{\mathrm{hist}}^{\backslash\mathrm{fit},E}\cdot\left|\mathbf{S}_{\mathrm{hist}}^{\backslash\mathrm{fit},F}\right|\cdot\left(1-\mathbf{P}_{\mathrm{loss}}^{\backslash\mathrm{fit}}\right)$
$P_{\mathrm{good}}^{\mathrm{min}}\leftarrow\mathrm{min}\left[S_{\mathrm{hist}}^{\mathrm{min}},\ \mathrm{min}\left(\mathbf{P}_{\mathrm{good}}^{\backslash\mathrm{NaN},>0}\right)\right]$
@@
$\mathbf{P}_{\mathrm{good}}\leftarrow 0\ \ \mathrm{if}\ S_{\mathrm{hist}}^{\backslash\mathrm{fit},r,E}\leq 0$
@@
$\mathbf{D}_{\mathrm{fit}}\leftarrow\mathbf{D}_{\mathrm{fit}}\cup\mathrm{random{\_}choice}\left(\mathbf{D}^{\backslash\mathrm{fit}},\ \mathbf{P}_{\mathrm{good}},\ N_{\mathrm{good}}\right)$
@@
The selection of the training structures to be fitted is given in Algorithm 1 and will be explained step-by-step in this paragraph. In each training step, the lADS algorithm utilizes a fraction of $p_{\mathrm{fit}}$ of all training structures that are still available for training, i.e., structures $r$ with an adaptive selection factor $S_{\mathrm{hist}}^{r,E}$ larger than zero. In general, these adaptive selection factors determine the training probability. They incorporate the representation quality history for the energy and forces of each structure. The starting value for each structure is $S_{\mathrm{hist}}^{r,E/F}=1$. From the $N_{\mathrm{fit}}$ structures, a fraction of $p_{\mathrm{good}}$ is selected from well represented structures, while the remaining $N_{\mathrm{bad}}$ structures are not yet well represented. $p_{\mathrm{good}}$ is initialized before the first step as zero and will be adapted as shown in Algorithm 2. By employing good and bad data, the stability–plasticity balance of retaining old expertise and integrating new knowledge can be improved. Moreover, this approach is required for lADS to sort out redundant, well represented data.
@@
To obtain the training probabilities, the maximum loss contributions for energies and forces $L_{\mathrm{old}}^{\mathrm{max},E/F}$ are determined among all previously calculated loss contributions of structures available for training. We note that the loss vectors $\mathbf{L}_{\mathrm{old}}^{E/F}$ are initialized before the first step as $\mathbf{NaN}$, i.e., a vector of not a number (NaN) entries. If there is at least one structure with a loss contribution greater than zero, the probability based on the loss $\mathbf{P}_{\mathrm{loss}}^{E/F}$ is calculated as the loss vector divided by the maximum loss contribution. Thereby, the minimum value is set to the mean loss contribution $\overline{L}_{\mathrm{old}}^{E/F}$ divided by the maximum loss contribution. The mean loss contributions for energy and forces are initialized before the first step as infinity. $\mathbf{P}_{\mathrm{loss}}^{E/F}$ increases the probability that badly represented structures are selected, while providing equal probabilities for well represented structures so that even very well represented structures have a chance of being selected. If no loss contribution is greater than zero, $\mathbf{P}_{\mathrm{loss}}^{E/F}=\mathbf{1}$ is applied. The probability vectors for energy and forces are multiplied and their square root is taken, resulting in $\mathbf{P}_{\mathrm{loss}}$. Subsequently, $\mathbf{P}_{\mathrm{loss}}$ is multiplied by the adaptive selection factor vectors of energies and forces. While $\mathbf{P}_{\mathrm{loss}}$ contains information about the current representation quality, the adaptive selections factors account for the history of the representation quality. The maximum value of the resulting vector $\mathbf{P}_{\mathrm{bad}}$ is determined to be $P_{\mathrm{bad}}^{\mathrm{max}}$ that has a lower bound of $S_{\mathrm{hist}}^{\mathrm{max}}$. The latter also defines the maximum value in $\mathbf{S}_{\mathrm{hist}}$. All $\mathrm{NaN}$ in $\mathbf{P}_{\mathrm{bad}}$ are set to $P_{\mathrm{bad}}^{\mathrm{max}}$, i.e., structures trained for the first time have the highest probability in $P_{\mathrm{bad}}$. For unavailable training structures, the respective vector entries are set to zero. The final probability is obtained by dividing the vector by the sum of its entries. Afterwards, $N_{\mathrm{bad}}$ structures are randomly selected from the training data $D$, with the respective probabilities $\mathbf{P}_{\mathrm{bad}}$.
@@
To choose the $N_{\mathrm{good}}$ structures, $1-\mathbf{P}_{\mathrm{loss}}$ is multiplied by the adaptive selection factor vectors for those structures that were not yet selected, resulting in $\mathbf{P}_{\mathrm{good}}$. Therefore, high loss contributions lead to a lower probability in this case. The minimum value larger than 0 is determined as $P_{\mathrm{good}}^{\mathrm{min}}$ that has an upper bound of $S_{\mathrm{hist}}^{\mathrm{min}}$. The latter also defines the minimum value of the adaptive selection factor of the structures available for training. $P_{\mathrm{good}}^{\mathrm{min}}$ will be employed in $\mathbf{P}_{\mathrm{good}}$ if the vector entry is $\mathrm{NaN}$ or the respective $P_{\mathrm{loss}}^{r}$ is $1$, i.e., structures trained for the first time and those with the highest loss have the lowest probability in $P_{\mathrm{good}}$. Analogously to $\mathbf{P}_{\mathrm{bad}}$, $\mathbf{P}_{\mathrm{good}}$ is set to zero for unavailable structures, and it is divided by the sum of its entries. Finally, the resulting random training data choice using the probabilities $\mathbf{P}_{\mathrm{good}}$ is combined with the data selected using $\mathbf{P}_{\mathrm{bad}}$.
@@
$\quad\quad\overline{L}_{Ti}^{E/F}\leftarrow\mathrm{mean}\left(\mathbf{L}_{Ti}^{\backslash\mathrm{NaN},S_{\mathrm{hist}}^{r,E}\geq-1,E/F}\right)$
@@
[…]
@@
+$\quad\quad\quad\ \ \ +\left\{\mathrm{max}\left[0,\ \left(S_{\mathrm{hist}}^{r,F}\right)^{-1}-1\right]\right\}^{2}>\left[\left(S_{\mathrm{hist}}^{\mathrm{max}}\right)^{-1}-1\right]^{2}$
@@
$\overline{L}_{\mathrm{new}}^{E/F}\leftarrow\mathrm{mean}\left(\mathbf{L}_{\mathrm{old}}^{\backslash\mathrm{NaN},S_{\mathrm{hist}}^{r,E}\geq-1,E/F}\right)$
@@
To update the selection-determining properties (Algorithm 2), the loss contributions $\mathbf{L}_{\mathrm{new}}^{E/F}$ of the currently chosen structures are calculated. In addition, loss contributions $\mathbf{L}_{Ti}^{E/F}$ are calculated for which the deviation is scaled by one of four threshold factors $T_{i}$, with $i={1,2,3,4}$. These vectors are initialized before the first step as $\mathbf{NaN}$. For each of these vectors the mean $\overline{L}_{Ti}^{E/F}$ is determined for available structures and those assigned to be redundant.
@@
If the new energy or force loss contribution of a structure $r$ is larger than the mean loss contribution applying the largest threshold factor $\overline{L}_{T4}^{E/F}$, then the exclusion strike counter $X^{r}$ of that structure will be increased by one. The starting value of this counter is zero for each structure. If the loss contributions for energy and forces are lower than the threshold, then the counter will be reset to zero. Afterwards, the adaptive selection factors $S_{\mathrm{hist}}^{r,E/F}$ are individually updated for energies and forces. If the new loss contribution ${L}_{\mathrm{new}}^{r,E/F}$ is not lower than the first threshold $\overline{L}_{T1}^{E/F}$, $S_{\mathrm{hist}}^{r,E/F}$ will have a lower bound of one. An upper bound of one will be applied if the value is not higher than the second threshold $\overline{L}_{T2}^{E/F}$. In this way, only well/bad represented structures can get a small/large $S_{\mathrm{hist}}^{r,E/F}$ value, while the $S_{\mathrm{hist}}^{r,E/F}$ values of structures with intermediate representation quality retain around one. Subsequently, $S_{\mathrm{hist}}^{r,E/F}$ is modified by large and small decrease factors $F_{--}$ and $F_{-}$ as well as small and large increase factors $F_{+}$ and $F_{++}$ depending on the value of the new loss contribution ${L}_{\mathrm{new}}^{r,E/F}$ compared to the threshold values $\overline{L}_{Ti}^{E/F}$ and the old loss contribution ${L}_{\mathrm{old}}^{r,E/F}$. These decrease and increase factors are calculated from the hyperparameters $N_{-/--}$ and $N_{+/++}$ by $F_{-/--}=(S_{\mathrm{min}})^{(N_{-/--})^{-1}}$ and $F_{+/++}=(S_{\mathrm{max}})^{(N_{+/++})^{-1}}$. Therefore, $N_{-/--}$ and $N_{+/++}$ define the number of repeated applications of the respective factor until $S_{\mathrm{min}}$ or $S_{\mathrm{max}}$ is reached. Exceeding these lower and upper bounds leads to the exclusion of the associated structure from training, as described below.

### v2

# Lifelong Machine Learning Potentials for Chemical Reaction Network Explorations

@@ Abstract
Recent developments in computational chemistry facilitate the automated quantum chemical exploration of chemical reaction networks for the in-silico prediction of synthesis pathways, yield, and selectivity. However, the underlying quantum chemical energy calculations require vast computational resources, limiting these explorations severely in practice. Machine learning potentials (MLPs) offer a solution to increase computational efficiency, while retaining the accuracy of reliable first-principles data used for their training. Unfortunately, MLPs will be limited in their generalization ability within chemical (reaction) space, if the underlying training data are not representative for a given application. Within the framework of automated reaction network exploration, where new reactants or reagents composed of any elements from the periodic table can be introduced, this lack of generalizability will be the rule rather than the exception. Here, we therefore evaluate the benefits of the lifelong MLP concept in this context. Lifelong MLPs push their adaptability by efficient continual learning of additional data. We propose an improved learning algorithm for lifelong adaptive data selection yielding efficient integration of new data while previous expertise is preserved. In this way, we can reach chemical accuracy in reaction search trials.
Lifelong Machine Learning Potentials, Chemical Reaction Networks, Continual Resilient (CoRe) Optimizer, Lifelong Adaptive Data Selection
@@ 1 Introduction
Accordingly, fast quantum chemical approaches are needed. However, they are plagued by drastic approximations that compromise not only energies but also molecular structures [32, 33]. As a consequence, a CRN constructed with an approximate quantum chemical approach (such as tight-binding density functional or semi-empirical theories) does not necessarily represent the network that would be obtained with a more reliable approach (such as density functional theory (DFT)). Hence, node and connection fidelity can be compromised. Local refinement by applying more accurate methods for the a-posteriori reoptimization of stable intermediates and transition states could improve molecular structures and, most importantly, their relative energies [34, 35, 36]. However, this requires a CRN to be qualitatively correct. Hence, the stationary points optimized with a fast approach must be complete as must be their connections by elementary reaction steps. It is therefore key to devise first-principles-based methods that are (I) fast to allow for efficient reactive event screening and that have (II) high structural fidelity so that the topology and structure of a CRN is not compromised. Hence, machine learning potentials (MLPs) hold the greatest promise to yield both, efficiency and structure fidelity [37, 38, 39, 40, 41, 42, 43, 44].
@@
However, the requirements for MLPs in CRN explorations are challenging. (I) They must be quickly initializable and (II) they must not come with an unbearable overhead for their training. It is precisely the goal of universal MLPs to deliver a parametrization that is as general as possible to overcome the initialization issue and to avoid training at runtime [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. To achieve this goal, models with universal structure representations are trained on very large and diverse reference data sets. Unfortunately, an out-of-the-box approach without further fine-tuning of the universal foundation model on some specialized data of the given chemical system can suffer from insufficient accuracy [56, 57]. In order to tackle this challenge, we therefore proposed the concept of lifelong machine learning potentials (lMLP) to alleviate the overhead of training [58]. In contrast to conventionally trained MLPs, lMLPs can continuously adapt to additional data without forgetting previous knowledge, while they do not require training again on all previous data. Hence, they can offer high accuracy for the parts of chemical space studied, but they may still require additional quantum chemical calculations during their application. We note that the lifelong learning process cannot only start from scratch but also from the parametrization of a pre-trained universal foundation model. In contrast to conventional fine-tuning of such a foundation model, lifelong learning can be continued for much more than one fine-tuning iteration.
@@
In this work, we assess the reliability, accuracy, and efficiency of lMLPs for CRN explorations. Section 2 summarizes the concept of lMLPs and introduces an improved algorithm for lifelong adaptive data selection. The computational details are compiled in Section 3. Section 4 starts with a presentation of the $\text{HCN}+\text{H}_{2}\text{O}$ CRN studied in this work. Afterwards, continual learning of an lMLP on the CRN data is analyzed and compared to conventional iterative learning and DFT reference data.
@@ 2.1 Lifelong Machine Learning Potentials
Figure 1: (a) In iterative learning, a new machine learning model is trained from scratch on all data to integrate additional training data. (b) In lifelong learning, the training of a model is continued employing added data and only a subset of previous data to prevent forgetting. In this way, previously acquired knowledge of the model is exploited. We note that lifelong learning is here only based on rehearsal of data, while it can also exploit model parameter regularization and the model architecture.
@@
With the help of continual learning, a closed and efficient loop in the evolution and application of an lMLP can be obtained (Figure 2). The process starts with some initial reference data. These data can be represented through universal descriptors such as element-embracing atom-centered symmetry functions (eeACSFs) [58], so that they can be learned by an lMLP utilizing, for example, the high-dimensional neural network potential (HDNNP) method [67, 68, 69] (see Supporting Information Section S1.1 of this work for an overview). Network expressivity by activation rank (NEAR) [70] can help with the choice of the neural network architecture. We note that the lMLP concept can also be employed with other methods, including approaches which combine descriptor and potential into a single learnable representation. Moreover, the initial parametrization of the MLP can be taken from a pre-trained universal foundation model instead of employing some random weight initialization. In any case, the representation must be able to handle any chemical structure to yield a generally applicable lMLP. For example, many MLP descriptors cannot efficiently represent a system with many different chemical elements because their descriptor vector size increases with an increasing number of different elements. As a consequence, the application of conventional atom-centered symmetry functions [71] is often restricted to systems with at most four different elements because of computational costs. By combination of (molecular) structure information with element information of the periodic table, eeACSFs can overcome this limitation [72, 58] (see Supporting Information Section S1.2 of this work for an overview). Moreover, in Supporting Information Section S1.2 of this work we propose alternative functions based on the bump function to represent the radial and angular structure within eeACSFs and we introduce the cube root-scaled-shifted (crss) scaling function for eeACSFs.
@@
To be able to deal with molecules with different total charges or spin multiplicities, individual lMLPs for each state can be trained. Alternatively, the second-generation HDNNP base model can be replaced by other models that can handle multiple states (and long-range interactions beyond the cutoff radius employed in the MLP descriptor), such as fourth-generation HDNNPs [73]. Moreover, continual learning can also be applied to train atomic-property machine learning models (e.g., for atomic charges and spins) to calculate these efficiently, e.g., during molecular dynamics simulations.
@@
Uncertainty quantification can be exploited in applications of the trained model to probe a pre-defined accuracy on the fly and to identify the need for new data points. In fact, uncertainty quantification is necessary to enable application of an lMLP at every training stage because it provides the confidence interval for analysis of the results. Uncertainties can be obtained, for example, by an ensemble or committee model [74, 75, 76, 77]. This can provide a quantitative estimate of small errors, while large errors can only be flagged. However, an indication of large errors is sufficient because only the need for additional data must be revealed. Those are then produced and fed into the continual training process until all errors are found to be sufficiently small.
@@ 2.2 Lifelong Adaptive Data Selection
Figure 3: Simplified scheme of the core concept behind lADS to continuously reduce and clean the data during training. (I) Data is redundant if it is seldom trained but still well represented. (II) Data is likely to be inconsistent if it is very often trained but still badly represented.
The main ideas behind lADS (Figure 3) are that, on the one hand, training data will be redundant if they are seldom trained in continual learning but still well represented. On the other hand, it is likely that data will be inconsistent with the majority of data if they are very often trained but still poorly represented. Consequently, lADS requires only a fraction of all training structures in each training step, applying a biased random selection. The probability for training of a given structure thereby needs to be adjusted according to how well the structure is represented by the model and how stable the quality of this representation is. In this way, both aforementioned ideas can be exploited. Moreover, learning can also focus on the insufficiently represented training data.
@@
with the loss functions $\mathcal{L}$ of the energy $E^{r}$ and the Cartesian components of the atomic forces $F_{\alpha,n}^{r}$ of the $N_{\mathrm{atom}}^{r}$ atoms $n$ of each structure $r$,
@@
The main difference of the lADS algorithm presented in this work compared to its original implementation in Reference [58] is (I) the separate consideration of energy and force loss contributions. (II) the advanced algorithm adjusts the number of fitted structures per training step. (III) it employs an improved scheme to calculate the training probabilities and (IV) to update the selection determining properties. (V) it includes a maximum number of structures that can be classified as redundant per step and (VI) adds a scheme for fast integration of additional training data. In addition, (VII) it enables removing loss gradient contributions of inconsistent training data in the history of the CoRe optimizer. The following four paragraphs will explain all details of the lADS algorithm.
@@
We use the following naming conventions: $\mathbf{D}$ is a vector of data points. $\mathbf{L}$ are loss contributions for each data point. $\mathbf{P}$ are probabilities for each data point to be selected for training. $\mathbf{S}_{\mathrm{hist}}$ are the adaptive selection factors employed in the probability determination of each data point. $N$ is a count of a quantity. Hyperparameters are defined in the main text.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The selection of the training structures to be fitted is given in Algorithm 1 and will be explained step-by-step in this paragraph. In each training step, the lADS algorithm utilizes a fraction of $p_{\mathrm{fit}}$ of all training structures that are still available for training, i.e., structures $r$ with an adaptive selection factor $S_{\mathrm{hist}}^{r,E}$ larger than zero. In general, these adaptive selection factors determine the training probability. They incorporate the representation quality history for the energy and forces of each structure. The starting value for each structure is $S_{\mathrm{hist}}^{r,E/F}=1$. From the $N_{\mathrm{fit}}$ structures, a fraction of $p_{\mathrm{good}}$ is selected from well represented structures, while the remaining $N_{\mathrm{bad}}$ structures are not yet well represented. $p_{\mathrm{good}}$ is initialized before the first step as zero and will be adapted as shown in Algorithm 2 in the next paragraph. By employing good and bad data, the stability–plasticity balance of retaining old expertise and integrating new knowledge can be improved. Moreover, training on good data is required for lADS to sort out redundant, well represented data.
@@
To obtain the training probabilities, the maximum loss contributions for energies and forces $L_{\mathrm{old}}^{\mathrm{max},E/F}$ are determined among all previously calculated loss contributions of structures available for training. We note that the loss vectors $\mathbf{L}_{\mathrm{old}}^{E/F}$ are initialized before the first step as $\mathbf{NaN}$, i.e., a vector of not a number (NaN) entries. If there is at least one structure with a loss contribution greater than zero, the probability based on the loss $\mathbf{P}_{\mathrm{loss}}^{E/F}$ is calculated as the loss vector divided by the maximum loss contribution. Thereby, the minimum value is set to the mean loss contribution $\overline{L}_{\mathrm{old}}^{E/F}$ divided by the maximum loss contribution. $\mathbf{P}_{\mathrm{loss}}^{E/F}$ increases the probability that badly represented structures are selected, while providing equal probabilities for well represented structures so that even very well represented structures have a chance of being selected. If no loss contribution is greater than zero, $\mathbf{P}_{\mathrm{loss}}^{E/F}=\mathbf{1}$ is applied. The probability vectors for energy and forces are multiplied and their square root is taken, resulting in $\mathbf{P}_{\mathrm{loss}}$. Subsequently, $\mathbf{P}_{\mathrm{loss}}$ is multiplied by the adaptive selection factor vectors of energies and forces. While $\mathbf{P}_{\mathrm{loss}}$ contains information about the current representation quality, the adaptive selections factors account for the history of the representation quality. The maximum value of the resulting vector $\mathbf{P}_{\mathrm{bad}}$ is determined to be $P_{\mathrm{bad}}^{\mathrm{max}}$ that has a lower bound of $S_{\mathrm{hist}}^{\mathrm{max}}$. The latter also defines the maximum value in $\mathbf{S}_{\mathrm{hist}}$. All $\mathrm{NaN}$ in $\mathbf{P}_{\mathrm{bad}}$ are set to $P_{\mathrm{bad}}^{\mathrm{max}}$, i.e., structures trained for the first time have the highest probability in $P_{\mathrm{bad}}$. For unavailable training structures, the respective vector entries are set to zero. The final probability is obtained by dividing the vector by the sum of its entries. Afterwards, $N_{\mathrm{bad}}$ structures are randomly selected from the training data $\mathbf{D}$, with the respective probabilities $\mathbf{P}_{\mathrm{bad}}$.
@@
To choose the $N_{\mathrm{good}}$ structures, $1-\mathbf{P}_{\mathrm{loss}}$ is multiplied by the adaptive selection factor vectors for those structures that were not yet selected, resulting in $\mathbf{P}_{\mathrm{good}}$. Therefore, high loss contributions lead to a lower probability in this case. The minimum value of all values larger than 0 is determined as $P_{\mathrm{good}}^{\mathrm{min}}$ that has an upper bound of $S_{\mathrm{hist}}^{\mathrm{min}}$. The latter also defines the minimum value of the adaptive selection factor of the structures available for training. $P_{\mathrm{good}}^{\mathrm{min}}$ will be employed in $\mathbf{P}_{\mathrm{good}}$ if the vector entry is $\mathrm{NaN}$ or the respective $P_{\mathrm{loss}}^{r}$ is $1$, i.e., structures trained for the first time and those with the highest loss have the lowest probability in $P_{\mathrm{good}}$. Analogously to $\mathbf{P}_{\mathrm{bad}}$, $\mathbf{P}_{\mathrm{good}}$ is set to zero for unavailable structures, and it is divided by the sum of its entries. Finally, the resulting random training data choice using the probabilities $\mathbf{P}_{\mathrm{good}}$ is combined with the data selected using $\mathbf{P}_{\mathrm{bad}}$. This combined data subsample $\mathbf{D}_{\mathrm{fit}}$ will be fitted in the training epoch.
@@
[…]
@@
$\mathbf{S}_{\mathrm{hist}}^{E/F}\leftarrow\mathrm{min}\left(S_{\mathrm{hist}}^{r,E/F},\ 1\right)\ \ \mathrm{if}\ L_{\mathrm{new}}^{r,E/F}\leq\overline{L}_{T2}^{E/F}$
@@
$\quad\quad\quad\ \ \ +\left\{\mathrm{max}\left[0,\ \left(S_{\mathrm{hist}}^{r,F}\right)^{-1}-1\right]\right\}^{2}>\left[\left(S_{\mathrm{hist}}^{\mathrm{max}}\right)^{-1}-1\right]^{2}$
@@
[…]
@@
To update the selection-determining properties (Algorithm 2), the loss contributions $\mathbf{L}_{\mathrm{new}}^{E/F}$ of the currently chosen structures are calculated. In addition, loss contributions $\mathbf{L}_{Ti}^{E/F}$ are calculated for which the deviation is scaled by one of four threshold factors $T_{i}$, with i=1,2,3,4i={1,2,3,4}. These vectors are initialized before the first step as $\mathbf{NaN}$. For each of these vectors the mean $\overline{L}_{Ti}^{E/F}$ is determined for the union of available and redundant structures.
@@
If the new energy or force loss contribution of a structure $r$ is larger than the mean loss contribution applying the largest threshold factor $\overline{L}_{T4}^{E/F}$, then the exclusion strike counter $X^{r}$ of that structure will be increased by one. The starting value of this counter is zero for each structure. If the loss contributions for energy and forces are lower than the threshold, then the counter will be reset to zero. Afterwards, the adaptive selection factors $S_{\mathrm{hist}}^{r,E/F}$ are individually updated for energies and forces. If the new loss contribution ${L}_{\mathrm{new}}^{r,E/F}$ is not lower than the first threshold $\overline{L}_{T1}^{E/F}$, $S_{\mathrm{hist}}^{r,E/F}$ will have a lower bound of one. An upper bound of one will be applied if the value is not higher than the second threshold $\overline{L}_{T2}^{E/F}$. In this way, only well/badly represented structures can get a small/large $S_{\mathrm{hist}}^{r,E/F}$ value, while the $S_{\mathrm{hist}}^{r,E/F}$ values of structures with intermediate representation quality retain around one. Subsequently, $S_{\mathrm{hist}}^{r,E/F}$ is modified by large and small decrease factors $F_{--}$ and $F_{-}$ as well as small and large increase factors $F_{+}$ and $F_{++}$ depending on the value of the new loss contribution ${L}_{\mathrm{new}}^{r,E/F}$ compared to the threshold values $\overline{L}_{Ti}^{E/F}$ and the old loss contribution ${L}_{\mathrm{old}}^{r,E/F}$. These decrease and increase factors are calculated from the hyperparameters $N_{-/--}$ and $N_{+/++}$ by $F_{-/--}=(S_{\mathrm{min}})^{(N_{-/--})^{-1}}$ and $F_{+/++}=(S_{\mathrm{max}})^{(N_{+/++})^{-1}}$. Therefore, $N_{-/--}$ and $N_{+/++}$ define the number of repeated applications of the respective factor until $S_{\mathrm{min}}$ or $S_{\mathrm{max}}$ is reached. Exceeding these lower and upper bounds leads to the exclusion of the associated structure from training, as described below.